OpenStudy (anonymous):

Rationalize the denominator 6sqrt of 5 / 5sqrt of 2

OpenStudy (anonymous):



OpenStudy (mimi_x3):

\[\large \frac{6\sqrt{5}}{5\sqrt{2}} *\frac{5\sqrt{2}}{5\sqrt{2}} =\frac{30\sqrt{10}}{25*2} = \frac{30\sqrt{10}}{50} =\frac{3}{5} \sqrt{10} \]

Directrix (directrix):